# High-dimensional Gaussian and bootstrap approximations for robust means

### v3

@@ 3.2 Bootstrap consistency
Fix $c\in(1,\sqrt{1.5})$, and let Assumption 2.1 be satisfied with $m>2$. Let $n>24$. If $\varepsilon_{n},\varepsilon_{n}^{\prime}\in(0,1/2)$, with $\varepsilon_{n}$ as in (5) and $\varepsilon_{n}^{\prime}$ as in (9), and $\tilde{Z}\sim\mathsf{N}_{d}(0,\tilde{\Sigma}_{n})$ conditionally on $\tilde{X}_{1},\ldots,\tilde{X}_{n}$, it holds with probability at least $1-\frac{24}{n}$ that
@@
$\leq\mathfrak{A}_{n}+C\mathinner{\Biggl(\log^{2}(d)\mathinner{\biggl[\overline{\eta}_{n}^{1-\frac{2}{m}}+\mathinner{\Bigl(\frac{\log(dn)}{n}\Bigr)}^{\frac{1}{2}-\frac{1}{m}}\biggr]}\Biggr)}^{1/2},$
where $\mathfrak{A}_{n}$ is the upper bound on $\rho_{n,W}$ in (6) of Theorem 2.1 and $C$ is a constant depending only on $b_{1},b_{2},c$ and $m$.
@@
$S_{n,\text{MB}}=\frac{1}{\sqrt{N}}\sum_{i=1}^{N}\xi_{i}V_{i}\qquad\text{where}\qquad V_{i}=\mathinner{\bigl(\phi_{\hat{a}_{1},\hat{b}_{1}}(\tilde{Y}_{i,1}),\ldots,\phi_{\hat{a}_{d},\hat{b}_{d}}(\tilde{Y}_{i,d})\bigr)}^{\prime}\in\mathbb{R}^{d},\ i=1,\ldots,N.$
@@
Remark 3.1 ($m\geq 4$).
The results so far in this section have imposed $m>2$ only. In case $m\geq 4$, implying that the $X_{1,j}X_{1,k}$ have second moments for all $1\leq j,k\leq d$, one can choose
$\varepsilon_{n}^{\prime}=2\lambda_{1,c}\overline{\eta}_{n}+\lambda_{2,c}(N,d^{2})\frac{\log(d^{2}N)}{N},\qquad\text{for }c\in(1,\sqrt{1.5})$ (12)
where the second summand tends to zero faster in $N=n/2$ than the second summand in (9), cf. Remark D.2 in the appendix. The resulting estimator, $\Sigma_{n}^{\prime}$, say, satisfies
$\mathbb{P}\mathinner{\Biggl(\max_{1\leq j,k\leq d}\mathinner{\!\bigl\lvert\check{\Sigma}_{n,j,k}-\Sigma_{j,k}\bigr\rvert}>C\mathinner{\biggl[\overline{\eta}_{n}^{1-\frac{2}{m}}+\mathinner{\Bigl(\frac{\log(dn)}{n}\Bigr)}^{\frac{1}{2}}+\mathinner{\Bigl(\frac{\log(dn)}{n}\Bigr)}^{1-\frac{2}{m}}\biggr]}\Biggr)}\leq\frac{24}{n},$
where $C$ is a constant depending only on $b_{2},c$, and $m$. The bootstrap consistency of Theorem 3.2 also carries over with this improved rate when $\check{\Sigma}_{n}$ replaces $\tilde{\Sigma}_{n}$.
@@ 4 Normalized winsorized means
In practice one often normalizes the data by an estimate of $\sigma_{2,j}$ to bring the variables on the same scale. We now describe how the Gaussian and bootstrap approximations established so far remain valid upon normalization by the diagonal elements $\tilde{\sigma}_{n,j}=\tilde{\Sigma}_{n,j,j}^{1/2}$ of the robust estimator $\tilde{\Sigma}_{n}$ of $\Sigma$, cf. (10). Writing $D=\text{diag}(\sigma_{2,1},\ldots,\sigma_{2,d})$ and $\Sigma_{0}=D^{-1}\Sigma D^{-1}$ for the correlation matrix of the (centered) uncontaminated $X_{1},\ldots,X_{n}$, one has the following Gaussian approximation for the vector $S_{n,W,S}$ of normalized winsorized means with elements (we leave the quotients undefined if one of the variance estimators equals 0)
@@
Fix $c\in(1,\sqrt{1.5})$, and let Assumption 2.1 be satisfied with $m>2$. If $\varepsilon_{n},\varepsilon_{n}^{\prime}\in(0,1/2)$, with $\varepsilon_{n}$ as in (5) and $\varepsilon_{n}^{\prime}$ as in (9), then for $Z^{\prime}\sim\mathsf{N}_{d}(0,\Sigma_{0})$,
@@
$\leq C\mathinner{\biggl(\mathfrak{A}_{n}+\sqrt{\log(d)\log(dn)}\mathinner{\Bigl[\overline{\eta}_{n}^{1-\frac{2}{m}}+\mathinner{\Bigl(\frac{\log(dn)}{n}\Bigr)}^{\frac{1}{2}-\frac{1}{m}}\Bigr]}\biggr)},$ (14)
where $\mathfrak{A}_{n}$ is the upper bound on $\rho_{n,W}$ in (6) of Theorem 2.1 and $C$ is a constant depending only on $b_{1},b_{2},c$ and $m$.
@@
Fix $c\in(1,\sqrt{1.5})$, and let Assumption 2.1 be satisfied with $m>2$. Let $n>24$. If $\varepsilon_{n},\varepsilon_{n}^{\prime}\in(0,1/2)$, with $\varepsilon_{n}$ as in (5) and $\varepsilon_{n}^{\prime}$ as in (9), and $\tilde{Z}^{\prime}\sim\mathsf{N}_{d}(0,\tilde{\Sigma}_{n,0})$ conditionally on $\tilde{X}_{1},\ldots,\tilde{X}_{n}$, it holds with probability at least $1-\frac{24}{n}$ that
@@
$\leq\mathfrak{B}_{n}+C\mathinner{\Biggl(\log^{2}(d)\mathinner{\biggl[\overline{\eta}_{n}^{1-\frac{2}{m}}+\mathinner{\Bigl(\frac{\log(dn)}{n}\Bigr)}^{\frac{1}{2}-\frac{1}{m}}\biggr]}\Biggr)}^{1/2},$ (15)
where $\mathfrak{B}_{n}$ is the upper bound on $\rho_{n,W,S}$ in (4.1) of Theorem 4.1 and $C$ is a constant depending only on $b_{1},b_{2},c$ and $m$.
@@ 5 Trimmed means
Analogously to the winsorized means in (3), the trimmed mean in (16) can be implemented in a fully data-driven way: Its amount of trimming is governed by $\varepsilon_{n}$ in (5), which does not depend on any unknown population quantities. [^7] As is typical in the literature on inference under data contamination, an upper bound $\overline{\eta}_{n}$ on the contamination rate must be supplied, however. This sets this trimmed mean apart from the one studied in Theorem 2 in Resende (2024) where the amount of trimming depends on $m$. Furthermore, the bound in Theorem 5.1 below is valid over the larger family of sets $\mathcal{H}$. In analogy to $\rho_{n,W}$ define
@@
Fix $c\in(1,\sqrt{1.5})$, and let Assumption 2.1 be satisfied with $m>2$. If $\varepsilon_{n}\in(0,1/2)$, with $\varepsilon_{n}$ as in (5), then
@@
where $\mathfrak{A}_{n}$ is the upper bound on $\rho_{n,W}$ in (6) of Theorem 2.1 and $C$ is a constant depending only on $b_{1},b_{2},c$ and $m$.
@@
Theorem 5.1 is proven by establishing that $S_{n,T}$ is sufficiently close to $S_{n,W}$ in the supremum-norm order to use a Gaussian anti-concentration inequality to deduce (17) from (6) in Theorem 2.1. This explains the presence of the additional summand on the right-hand side in (17). The same discussion as that following Theorem 2.1 also applies to the trimmed mean in (16).
@@
Fix $c\in(1,\sqrt{1.5})$, and let Assumption 2.1 be satisfied with $m>2$. Let $n>24$. If $\varepsilon_{n},\varepsilon_{n}^{\prime}\in(0,1/2)$, with $\varepsilon_{n}$ as in (5) and $\varepsilon_{n}^{\prime}$ as in (9), and $\tilde{Z}\sim\mathsf{N}_{d}(0,\tilde{\Sigma}_{n})$ conditionally on $\tilde{X}_{1},\ldots,\tilde{X}_{n}$, it holds with probability at least $1-\frac{24}{n}$ that
@@
$\leq\mathfrak{C}_{n}+C\mathinner{\Biggl(\log^{2}(d)\mathinner{\biggl[\overline{\eta}_{n}^{1-\frac{2}{m}}+\mathinner{\Bigl(\frac{\log(dn)}{n}\Bigr)}^{\frac{1}{2}-\frac{1}{m}}\biggr]}\Biggr)}^{1/2},$
where $\mathfrak{C}_{n}$ is the upper bound on $\rho_{n,T}$ in (17) of Theorem 5.2 and $C$ is a constant depending only on $b_{1},b_{2},c$ and $m$.
@@ Appendix A A useful decomposition
We begin by outlining a decomposition of $S_{n,W}$ used in the proof of Theorem 2.1. For $p\in(0,1)$ and a random variable $X$, denote by $Q_{p}(X)$ the $p$-quantile of the distribution of $X$, that is
$Q_{p}(Z)=\inf\mathinner{\bigl\{z\in\mathbb{R}\mathrel{\mathop{\ordinarycolon}}\mathbb{P}(X\leq z)\geq p\bigr\}}.$
Suppressing the dependence on $c\in(1,\sqrt{1.5})$, which is fixed throughout, for $\varepsilon_{n}\in(0,0.5)$ let
$\underline{\alpha}_{j}\mathrel{\mathop{\ordinarycolon}}=\underline{\alpha}_{c,j}\mathrel{\mathop{\ordinarycolon}}=Q_{\varepsilon_{n}-c^{-1}\varepsilon_{n}}(X_{1,j}),\qquad\text{and}\qquad\overline{\alpha}_{j}\mathrel{\mathop{\ordinarycolon}}=\overline{\alpha}_{c,j}\mathrel{\mathop{\ordinarycolon}}=Q_{\varepsilon_{n}+c^{-1}\varepsilon_{n}}(X_{1,j})$ (A.1)
@@
$\underline{\beta}_{j}\mathrel{\mathop{\ordinarycolon}}=\underline{\beta}_{c,j}\mathrel{\mathop{\ordinarycolon}}=Q_{1-\varepsilon_{n}-c^{-1}\varepsilon_{n}}(X_{1,j}),\qquad\text{and}\qquad\overline{\beta}_{j}\mathrel{\mathop{\ordinarycolon}}=\overline{\beta}_{c,j}\mathrel{\mathop{\ordinarycolon}}=Q_{1-\varepsilon_{n}+c^{-1}\varepsilon_{n}}(X_{1,j}).$ (A.2)
By definition $\hat{\alpha}_{j}=\tilde{X}_{\lceil\varepsilon_{n}n\rceil,j}^{*}$ and $\hat{\beta}_{j}=\tilde{X}_{\lceil(1-\varepsilon_{n})n\rceil,j}^{*}$. Lemma G.2 and the union bound, together with obvious monotonicity properties of $(a,b)\mapsto\phi_{a,b}$, show that with probability at least $1-\frac{4}{n}$ we have simultaneously for $j=1,\ldots,d$ (the inequalities $\phi_{\underline{\alpha}_{j},\underline{\beta}_{j}}\leq\phi_{\hat{\alpha}_{j},\hat{\beta}_{j}}\leq\phi_{\overline{\alpha}_{j},\overline{\beta}_{j}}$ and hence)
@@
where here and in the sequel (i) all inequalities between vectors are understood elementwise, and (ii) with some abuse of notation we define, for a vector $x\in\overline{\mathbb{R}}^{m}$, say, and a real number $a$, the sum $x+a$ coordinatewise as $(x_{1}+a,\ldots,x_{m}+a)^{\prime}$ , with the usual convention that $\infty+a=\infty$ and $\infty+a=-\infty$. In the next section we study the left- and right-hand sides of the previous display.

### v4

@@ 3.2 Bootstrap consistency
Let Assumption 2.1 be satisfied with $m>2$. Let $n>24$. If $\varepsilon_{n},\varepsilon_{n}^{\prime}\in(0,1/2)$, with $\varepsilon_{n}$ as in (4) and $\varepsilon_{n}^{\prime}$ as in (7), and $\tilde{Z}\sim\mathsf{N}_{d}(0,\tilde{\Sigma}_{n})$ conditionally on $\tilde{X}_{1},\ldots,\tilde{X}_{n}$, it holds with probability at least $1-\frac{24}{n}$ that
@@
$\leq\mathfrak{A}_{n}+C\mathinner{\Biggl(\log^{2}(d)\mathinner{\biggl[\overline{\eta}_{n}^{1-\frac{2}{m}}+\mathinner{\Bigl(\frac{\log(dn)}{n}\Bigr)}^{1-\frac{1}{(m/2)\wedge 2}}\biggr]}\Biggr)}^{1/2},$
where $\mathfrak{A}_{n}$ is the upper bound on $\rho_{n,W}$ in (5) of Theorem 2.1 and $C$ is a constant depending only on $b_{1},b_{2},\lambda_{1}^{\prime},\lambda_{2}^{\prime}$ and $m$.
@@
$S_{n,\text{MB}}=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\xi_{i}V_{i}\qquad\text{where}\qquad V_{i}=\mathinner{\bigl(\phi_{\hat{a}_{1},\hat{b}_{1}}(\tilde{X}_{i,1})-\tilde{\mu}_{n,1},\ldots,\phi_{\hat{a}_{d},\hat{b}_{d}}(\tilde{X}_{i,d})-\tilde{\mu}_{n,d}\bigr)}^{\prime}\in\mathbb{R}^{d}.$
@@ 4 Normalized winsorized means
In practice one often normalizes the data by estimates of the $\sigma_{2,j}$ to bring the variables on the same scale. We now describe how the Gaussian and bootstrap approximations established so far remain valid upon normalization by the diagonal elements $\tilde{\sigma}_{n,j}=\tilde{\Sigma}_{n,j,j}^{1/2}$ of the robust estimator $\tilde{\Sigma}_{n}$ of $\Sigma$, cf. (8). Writing $D=\text{diag}(\sigma_{2,1},\ldots,\sigma_{2,d})$ and $\Sigma_{0}=D^{-1}\Sigma D^{-1}$ for the correlation matrix of the (centered) uncontaminated $X_{1},\ldots,X_{n}$, one has the following Gaussian approximation for the vector $S_{n,W,S}$ of normalized winsorized means with elements (we leave the quotients undefined if one of the variance estimators equals 0)
@@
Let Assumption 2.1 be satisfied with $m>2$. If $\varepsilon_{n},\varepsilon_{n}^{\prime}\in(0,1/2)$, with $\varepsilon_{n}$ as in (4) and $\varepsilon_{n}^{\prime}$ as in (7), then for $Z^{\prime}\sim\mathsf{N}_{d}(0,\Sigma_{0})$,
@@
$\leq C\mathinner{\Biggl(\mathfrak{A}_{n}+\sqrt{\log(d)\log(dn)}\mathinner{\biggl[\overline{\eta}_{n}^{1-\frac{2}{m}}+\mathinner{\Bigl(\frac{\log(dn)}{n}\Bigr)}^{1-\frac{1}{(m/2)\wedge 2}}\biggr]}\Biggr)},$ (12)
where $\mathfrak{A}_{n}$ is the upper bound on $\rho_{n,W}$ in (5) of Theorem 2.1 and $C$ is a constant depending only on $b_{1},b_{2},\lambda_{1},\lambda_{2},\lambda_{1}^{\prime},\lambda_{2}^{\prime}$ and $m$.
@@
Let Assumption 2.1 be satisfied with $m>2$. Let $n>24$. If $\varepsilon_{n},\varepsilon_{n}^{\prime}\in(0,1/2)$, with $\varepsilon_{n}$ as in (4) and $\varepsilon_{n}^{\prime}$ as in (7), and $\tilde{Z}^{\prime}\sim\mathsf{N}_{d}(0,\tilde{\Sigma}_{n,0})$ conditionally on $\tilde{X}_{1},\ldots,\tilde{X}_{n}$, it holds with probability at least $1-\frac{24}{n}$ that
@@
$\leq\mathfrak{B}_{n}+C\mathinner{\Biggl(\log^{2}(d)\mathinner{\biggl[\overline{\eta}_{n}^{1-\frac{2}{m}}+\mathinner{\Bigl(\frac{\log(dn)}{n}\Bigr)}^{1-\frac{1}{(m/2)\wedge 2}}\biggr]}\Biggr)}^{1/2},$ (13)
where $\mathfrak{B}_{n}$ is the upper bound on $\rho_{n,W,S}$ in (4.1) of Theorem 4.1 and $C$ is a constant depending only on $b_{1},b_{2},\lambda_{1}^{\prime},\lambda_{2}^{\prime}$ and $m$.
@@ 5 Trimmed means
Analogously to the winsorized means in (3), the trimmed mean in (14) can be implemented in a fully data-driven way: Its amount of trimming is governed by $\varepsilon_{n}$ in (4), which does not depend on any unknown population quantities. [^5] As is typical in the literature on inference under data contamination, an upper bound $\overline{\eta}_{n}$ on the contamination rate must be supplied, however. This sets this trimmed mean apart from the one studied in Theorem 2 in Resende (2024), where the amount of trimming depends on $m$. Furthermore, the bound in Theorem 5.1 below is valid over the larger family of sets $\mathcal{H}$ instead of $\mathcal{R}$. In analogy to $\rho_{n,W}$ define
@@
Let Assumption 2.1 be satisfied with $m>2$. If $\varepsilon_{n}\in(0,1/2)$, with $\varepsilon_{n}$ as in (4), then
@@
where $\mathfrak{A}_{n}$ is the upper bound on $\rho_{n,W}$ in (5) of Theorem 2.1 and $C$ is a constant depending only on $b_{1},b_{2},\lambda_{1},\lambda_{2}$ and $m$.
@@
Theorem 5.1 is proven by establishing that $S_{n,T}$ is sufficiently close to $S_{n,W}$ in the supremum-norm, in order to use a Gaussian anti-concentration inequality to deduce (15) from (5) in Theorem 2.1. This explains the presence of the additional summand on the right-hand side in (15). The same discussion as that following Theorem 2.1 also applies to the trimmed mean in (14).
@@
Let Assumption 2.1 be satisfied with $m>2$. Let $n>24$. If $\varepsilon_{n},\varepsilon_{n}^{\prime}\in(0,1/2)$, with $\varepsilon_{n}$ as in (4) and $\varepsilon_{n}^{\prime}$ as in (7), and $\tilde{Z}\sim\mathsf{N}_{d}(0,\tilde{\Sigma}_{n})$ conditionally on $\tilde{X}_{1},\ldots,\tilde{X}_{n}$, it holds with probability at least $1-\frac{24}{n}$ that
@@
$\leq\mathfrak{C}_{n}+C\mathinner{\Biggl(\log^{2}(d)\mathinner{\biggl[\overline{\eta}_{n}^{1-\frac{2}{m}}+\mathinner{\Bigl(\frac{\log(dn)}{n}\Bigr)}^{1-\frac{1}{(m/2)\wedge 2}}\biggr]}\Biggr)}^{1/2},$
where $\mathfrak{C}_{n}$ is the upper bound on $\rho_{n,T}$ in (15) of Theorem 5.1 and $C$ is a constant depending only on $b_{1},b_{2},\lambda_{1}^{\prime},\lambda_{2}^{\prime}$ and $m$.
In particular, $\tilde{\rho}_{n,T}\to 0$ in probability if $\sqrt{n\log(d)}\overline{\eta}_{n}^{1-\frac{1}{m}}\to 0$ and $\log(d)/n^{\frac{m-2}{5m-2}}\to 0$.
@@
6 Numerical results
The P-P (probability-probability) plots in Figures 2 and 3 below illustrate the theoretical results of the paper. To allow a comparison to existing Gaussian approximations to the distribution of $S_{n}=n^{-1/2}\sum_{i=1}^{n}(X_{i}-\mu)$, which all are for uncontaminated $X_{i}$, we set $\overline{\eta}_{n}=0$ throughout. The $X_{i,j}$ are i.i.d., following a $t(\nu)$-distribution with $\nu$ degrees of freedom such that $\mu=0$ and $\Sigma=\text{diag}(\nu/(\nu-2),\ldots,\nu/(\nu-2))$. We consider $\nu=3.01$ (Figure 2) and $\nu=4.01$ (Figure 3), corresponding to the $X_{i,j}$ having all moments strictly less than $3.01$ and $4.01$, respectively. Throughout, $n=200$ and we study $d\in\mathinner{\{500,5{,}000\}}$. When implementing $S_{n,W}$ and $S_{n,T}$$\lambda_{1}$ and $\lambda_{1}^{\prime}$ are irrelevant as $\overline{\eta}_{n}=0$. We set $\lambda_{2}=0.1$ throughout. We implement $\tilde{\Sigma}_{n}$ with $\lambda_{2}^{\prime}=0.07$ since then the number of observations winsorized in implementing $\tilde{\Sigma}_{n}$ is typically the same as in implementing $S_{n,W}$ and $S_{n,T}$, i.e., $\varepsilon_{n}n=\lambda_{2}\log(dn)\approx\lambda_{2}^{\prime}\log(d^{2}n)=\varepsilon_{n}^{\prime}n$, which is found to work well. All results are based on 10,00010{,}000 replications and we write $x||_{\infty}=\max_{1\leq j\leq d}|x_{j}|$ for $x\in\mathbb{R}^{d}$.
[…]
Figure 3 increases $\nu$ to $4.01$ from $3.01$. As expected, this generally improves the performance of all approximations. In conclusion, the Gaussian and bootstrap approximations to the winsorized mean work well in all settings considered. We note, however, that whenever $n\varepsilon_{n}$ or $n\varepsilon_{n}^{\prime}$ cross an integer value, the number of winsorized observations changes, which can lead to a “discontinuity” in the performance. We also emphasize again that we have investigated scenarios without contamination. The normal and bootstrap approximations for the sample mean break down completely under contamination.
$\nu=3.01$$n=200$ and $d=500$
$\nu=3.01$$n=200$ and d=5,000d=5{,}000
[…]
$\nu=4.01$$n=200$ and $d=500$
$\nu=4.01$$n=200$ and d=5,000d=5{,}000
[…]
@@ Appendix A A useful decomposition
We begin by outlining a decomposition of $S_{n,W}$ used in the proof of Theorem 2.1 along with some further notation used in the appendix. For $p\in(0,1)$ and a random variable $X$, denote by $Q_{p}(X)$ the $p$-quantile of the distribution of $X$, that is
$Q_{p}(X)=\inf\mathinner{\bigl\{z\in\mathbb{R}\mathrel{\mathop{\ordinarycolon}}\mathbb{P}(X\leq z)\geq p\bigr\}}.$
For $A_{+}=1-\lambda_{1}^{-1}\mathds{1}\mathinner{\{\overline{\eta}_{n}>0\}}$ and $A_{-}=1+\lambda_{1}^{-1}\mathds{1}\mathinner{\{\overline{\eta}_{n}>0\}}$, define (recall $\varepsilon_{n}$ from (4))
$c_{1,n}\mathrel{\mathop{\ordinarycolon}}=-A_{+}W_{0}\mathinner{\bigl(-e^{-(\frac{\log(dn)}{\varepsilon_{n}n}+A_{+})/A_{+}}\bigr)}\in(0,A_{+})$ (A.1)
@@
$c_{2,n}\mathrel{\mathop{\ordinarycolon}}=-A_{-}W_{-1}\mathinner{\bigl(-e^{-(\frac{\log(dn)}{\varepsilon_{n}n}+A_{-})/A_{-}}\bigr)}\in(A_{-},\infty);$ (A.2)
here $W_{0}$ and $W_{-1}$ denote the principal and lower branch, respectively, of Lambert’s $W$ function (cf., e.g., Corless et al. (1996)). Because $A_{+}\leq A_{-}$, note that $0<c_{1,n}<c_{2,n}<\infty$. Next, for $j=1,\ldots,d$, let
$\underline{\alpha}_{j}\mathrel{\mathop{\ordinarycolon}}=Q_{c_{1,n}\varepsilon_{n}}(X_{1,j})\qquad\text{and}\qquad\overline{\alpha}_{j}\mathrel{\mathop{\ordinarycolon}}=Q_{c_{2,n}\varepsilon_{n}}(X_{1,j})$ (A.3)
as well as
$\underline{\beta}_{j}\mathrel{\mathop{\ordinarycolon}}=Q_{1-c_{2,n}\varepsilon_{n}}(X_{1,j})\qquad\text{and}\qquad\overline{\beta}_{j}\mathrel{\mathop{\ordinarycolon}}=Q_{1-c_{1,n}\varepsilon_{n}}(X_{1,j}),$ (A.4)
In the appendix only we shall sometimes impose
$\mathfrak{T}_{n}^{(1)}\mathrel{\mathop{\ordinarycolon}}=2\varepsilon_{n}+\frac{\log(dn)}{n}+\sqrt{\mathinner{\Bigl(\frac{\log(dn)}{n}\Bigr)}^{2}+4\frac{\log(dn)}{n}\varepsilon_{n}}<\frac{1}{2},$ (A.5)
which, however, is not imposed for any of our Gaussian and bootstrap approximations in this paper, as these will be argued to hold trivially when (A.5) is not satisfied. By Lemma G.2 one has that (A.5) ensures that
$0<\varepsilon_{n}\min(c_{1,n},c_{2,n})\leq\varepsilon_{n}(c_{1,n}+c_{2,n})<1/2,$
implying that $\underline{\alpha}_{j}\leq\underline{\beta}_{j}$ and $\overline{\alpha}_{j}\leq\overline{\beta}_{j}$, such that the decompositions in (A.7) and (A.8) below are well-defined.
By definition, $\hat{\alpha}_{j}=\tilde{X}_{\lceil\varepsilon_{n}n\rceil,j}^{*}$ and $\hat{\beta}_{j}=\tilde{X}_{\lceil(1-\varepsilon_{n})n\rceil,j}^{*}$. Lemma G.3 and the union bound, together with obvious monotonicity properties of $(a,b)\mapsto\phi_{a,b}$, show that with probability at least $1-\frac{4}{n}$ we have simultaneously for $j=1,\ldots,d$ (the inequalities $\phi_{\underline{\alpha}_{j},\underline{\beta}_{j}}\leq\phi_{\hat{\alpha}_{j},\hat{\beta}_{j}}\leq\phi_{\overline{\alpha}_{j},\overline{\beta}_{j}}$ and hence)
@@
where here and in the sequel (i) all inequalities between vectors are understood elementwise, and (ii) with some abuse of notation we define, for a vector $x\in\overline{\mathbb{R}}^{m}$, say, and a real number $a$, the sum $x+a$ coordinate-wise as $(x_{1}+a,\ldots,x_{m}+a)^{\prime}$ , with the usual convention that $\infty+a=\infty$ and $\infty+a=-\infty$. In the next section we study the left- and right-hand sides of the previous display.